# Is this chart lying to me? Automating the detection of misleading visualizations

### v1

@@ Abstract
Misleading visualizations are a potent driver of misinformation on social media and the web. By violating chart design principles, they distort data and lead readers to draw inaccurate conclusions. Prior work has shown that both humans and multimodal large language models (MLLMs) are frequently deceived by such visualizations. Automatically detecting misleading visualizations and identifying the specific design rules they violate could help protect readers and reduce the spread of misinformation. However, the training and evaluation of AI models has been limited by the absence of large, diverse, and openly available datasets. In this work, we introduce Misviz, a benchmark of 2,604 real-world visualizations annotated with 12 types of misleaders. To support model training, we also release Misviz-synth, a synthetic dataset of 81,814 visualizations generated using Matplotlib and based on real-world data tables. We perform a comprehensive evaluation on both datasets using state-of-the-art MLLMs, rule-based systems, and fine-tuned classifiers. Our results reveal that the task remains highly challenging. We release Misviz, Misviz-synth, and the accompanying code.111github.com/UKPLab/arxiv2025-misviz
@@ 1 Introduction
Misleading visualizations are charts that distort the underlying data, typically by violating design principles, leading readers to draw inaccurate conclusions (Tufte and Graves-Morris, 1983; Pandey et al., 2015; Lauer and O’Brien, 2020; McNutt et al., 2020; Lo et al., 2022; Lisnic et al., 2023; Lan and Liu, 2025). While many arise from unintentional design errors, misleading visualizations are also deliberately crafted by malicious actors to spread disinformation and manipulate public understanding, especially during crises such as the COVID-19 pandemic, where misleading charts circulated widely on social media (Correll and Heer, 2017; Tartaglione and de Wit, 2025; Lisnic et al., 2023). Prior work has shown that both humans (Pandey et al., 2014, 2015; O’Brien and Lauer, 2018; Yang et al., 2021; Ge et al., 2023; Rho et al., 2023) and MLLMs (Bendeck and Stasko, 2025; Chen et al., 2025; Pandey and Ottley, 2025; Tonglet et al., 2025) are deceived by such visualizations in question-answering tasks.
The deceptive features in these charts, or misleaders (Lisnic et al., 2023), often reside in subtle details easily missed by readers, such as axis tick intervals. Furthermore, misleaders are highly diverse: the latest taxonomies identify over 70 distinct types spanning a wide variety of chart types, including bar charts, pie charts, and maps (Lo et al., 2022; Lan and Liu, 2025). In some cases, more than one misleader affects the same visualization (Lo et al., 2022). Figure 1 shows 12 real-world examples.
@@
Automatically classifying whether a visualization is misleading, and identifying which misleaders affect it, if any, can enable timely warnings to chart designers and readers and help prevent the spread of misinformation. This task is framed as a multi-label classification problem. While early work relied on rule-based systems called linters (Hopkins et al., 2020; Fan et al., 2022), recent studies have explored the use of MLLMs (Lo and Qu, 2025; Alexander et al., 2024). However, these approaches were evaluated on distinct datasets, which are either small or closed-access, limiting comparability and hindering progress.
In this work, we introduce Misviz, a large, diverse, and open benchmark comprising 2,604 real-world visualizations spanning 12 types of misleaders. It reflects scenarios in which vulnerable readers encounter misleading visualizations online. In Misviz, 70% of the visualizations contain up to three misleaders, while the remainder are non-misleading. To support model training, we also release Misviz-synth, a synthetic dataset generated with Matplotlib (Hunter, 2007) using real-world data tables. Misviz-synth simulates scenarios where chart designers seek assistance in identifying misleaders unintentionally introduced in their visualizations. It includes not only the visualizations but also their data tables and axis metadata, enabling the training of chart de-rendering models.
We conduct extensive experiments with three approaches: (a) state-of-the-art MLLMs, (b) a new rule-based linter that inspect axis metadata for design rule violations, and (c) new classifiers that take the visualization alone or in combination with the axis metadata as input. For (b) and (c), we fine-tune DePlot (Liu et al., 2023) to extract axis metadata. Our experiments address the following research questions (RQs). RQ1: Which type of model performs best on real-world or synthetic instances? RQ2: Can models trained on synthetic instances generalize to real-world cases? RQ3: Can models fine-tuned for axis extraction on synthetic instances generalize to real-world cases?
Our results show that MLLMs perform best on real-world visualizations, while linters and fine-tuned classifiers outperform them on synthetic ones, benefiting from the availability of training data for both axis extraction and misleader detection. While the fine-tuned DePlot can extract axes from Misviz-synth, it does not generalize well to Misviz, reducing the performance of the linter and classifier.
In summary, our contributions are as follows: (1) we introduce Misviz and Misviz-synth, the first large scale open datasets for misleading visualization detection, (2) we propose a new linter and a new classification method that combines image and extracted axis metadata as input, (3) we conduct comprehensive evaluation and error analysis.
@@ 2 Related work
Dataset Instances Misleader types Chart types % non-misleading Open access Real-world Multi-label Axis, table
@@
Misvisfix (Das and Mueller, 2025) 450 74 > 5 20
Misviz-synth (ours) 81,814 11 3 22
@@
The first attempts to detect misleading visualizations relied on rule-based systems called linters (McNutt and Kindlmann, 2018; Hopkins et al., 2020; Chen et al., 2022). These linters assume the availability of the underlying data table or the chart code, which restricts their applicability to real-world scenarios. Fan et al. (2022) and Biselli et al. (2025) overcome these limitations by extracting tables with OCR tools from real-world visualizations before applying rule checks. However, the accuracy of these real-world linters depends heavily on the quality of the intermediate OCR step, which can vary widely (Biselli et al., 2025). Real-world linters have previously been evaluated in small-scale user studies, with human-in-the-loop to correct OCR errors.
Others have explored the potential of MLLMs for the task. Lo and Qu (2025) evaluated four MLLMs with different prompts on a dataset of 150 real-world visualizations, sourced from the corpus of Lo et al. (2022) for misleading cases. They found that detection accuracy decreased as more misleader types were included in the prompt. Alexander et al. (2024) focused on GPT-4 (OpenAI, 2023), using visualizations from the social media platform X (Lisnic et al., 2023). However, access to this dataset requires a paid API, and reproducibility is further hindered by the frequent removal of the posts from the platform. In a parallel work, Das and Mueller (2025) proposed a prompt with which SOTA MLLMs achieve high accuracy on a subset of the corpus of Lo et al. (2022).
@@
Table 1 compares prior datasets with Misviz and Misviz-synth. Misviz is over fifteen times larger than the dataset of Lo and Qu (2025). Unlike Alexander et al. (2024), it does not rely on paid APIs and ensures long-term access to all instances by archiving them on the Wayback Machine.222web.archive.org Misviz-synth is three to ten times larger than other synthetic datasets, and include several more misleader types (Arif et al., 2024; Maciborski et al., 2025). In contrast to other synthetic datasets, Misviz-synth provides axis metadata, which is necessary to fine-tune DePlot for axis extraction and answer our research questions.
@@ 3.1 Selected misleaders
Misviz covers 12 types of misleaders, selected from the 74 categories defined in the taxonomy of Lo et al. (2022), based on three key criteria. First, we excluded misleaders that are rarely observed in real-world scenarios. To determine this, we used misleader frequency statistics from the corpus of Lo et al. (2022) and discarded all categories with fewer than three instances. Second, we removed misleaders that are more likely to confuse than deceive the reader. As noted by Lo et al. (2022), the taxonomy includes both misleaders that distort the underlying data, the focus of this work, and others that may hinder readability or clarity without altering the interpretation of the data, such as missing titles or overplotting. Third, we excluded misleaders that require specific domain knowledge to identify. For example, using red to represent Democrats and blue for Republicans in a U.S. politics chart constitutes a violation of color conventions, but detecting this misleader requires familiarity with U.S. politics. Such misleaders requires domain expertise that is not achievable with crowdworkers.
We define each selected misleader briefly below (Lo et al., 2022; Ge et al., 2023; Lan and Liu, 2025). Each of them is represented with an example in Figure 1 and in Appendix A.
Misrepresentation: the value labels displayed do not match the size of their visual encodings, e.g., bars may be drawn disproportionate to the corresponding numerical value.
@@
Inconsistent tick intervals: the ticks in one axis are evenly spaced but their values are not, e.g., the tick values sequence is 10, 20, 40, 45.
@@ 4.1 Data collection
We collect real-world data tables from two primary sources: Wikipedia and Our World in Data (OWID).555ourworldindata.org For Wikipedia, we use the tables included in the TabFact question answering dataset (Wenhu et al., 2020). OWID is an open-access data platform that covers a wide range of domains, including health, education, and the economy.
@@ 4.2 Synthetic visualizations generation
Figure 2: The two step process to create the synthetic visualizations of Misviz-synth based on real-world data.
@@
In the first step, given a raw real-world data table, we assign one or more types to each column. There are 10 different column types, explained in Appendix G. Based on these types, we select valid column combinations, requiring at least one numerical column and a natural key for each row. This natural key can consist of one or a combination of categorical or temporal columns. If the natural key involves multiple columns, we fix all but one to a constant value. For instance, if the numerical column is Win ratio and the natural key includes Football club and Year, we set Year to 2021 to compare the win ratios of football clubs for that year. Chart titles are automatically generated using templates based on the selected columns. We then determine which chart types, among bar, line, and pie charts, are suitable for the data. For example, line charts are only compatible with temporal data, not categorical data. Each valid configuration produces a “prepared” data table.
@@
For bar and line charts, we automatically extract axis metadata. Since pie charts do not have axes, they are excluded from this step. Axis metadata is stored as a table with four columns. (1) Seq indicates the sequential order of the tick marks along an axis. Horizontal axes are read from left to right, while vertical axes are read from bottom to top. (2) Axis indicates the name of the axis, e.g., x, y1, or y2, in case of dual axis. (3) Label is the tick label. (4) Relative position is a normalized float indicating the tick mark’s position, with spacing expressed relative to the distance between the first two tick marks. Examples of axis metadata are provided in Appendix H.
To increase visual diversity, we apply random design variations such as font size, background color, and axis label position. The complete list of variations is given in Appendix I.
Each instance is saved along with its title, data table, and axis metadata if available.
@@ 4.3 Data statistics
Misviz-synth contains 81,184 visualizations, distributed into stratified train (93%), validation (2%), and test sets (5%). Misviz-synth focuses on the three most common chart types in Misviz: bar (46%), line (47%), and pie charts (7%). Due to Matplotlib’s limitations, it does not include maps, and therefore omits the discretized continuous variable misleader, which is specific to maps. 78% of the visualizations are misleading. Unlike Misviz, each visualization in Misviz-synth contains at most one chart type and one misleader, making it slightly less diverse than Misviz.
@@ 5.1 Baselines
Zero-shot MLLMs We evaluate the zero-shot capabilities on one run of the following commercial and open-weight MLLMs: GPT-4.1 and GPT-o3 (OpenAI, 2023), Gemini-1.5-Flash and Gemini-1.5-Pro (Gemini-Team, 2024), Qwen-2.5-VL in 7B, 32B, and 72B variants (Qwen-Team, 2025), and InternVL3 in 8B, 38B, and 78B variants (Zhu et al., 2025). These models were selected due to their strong performance on the ChartQA benchmark (Masry et al., 2022). The prompt includes the task description as well as the definitions of the misleaders. The prompt is provided in Appendix J.
Rule-based linter We design a linter that detects misleaders by applying rule-based checks to the axis metadata of a visualization. Each misleader corresponds to a specific rule. If a visualization passes all checks, it is classified as having no misleaders. We rely on axis metadata rather than the underlying data table because it enables detection of a broader range of misleaders: truncated axis, inverted axis, dual axis, inconsistent tick intervals, inconsistent binning size, and inappropriate item order. Only the latter two misleaders could also be detected from the data table alone. Detailed descriptions of all rule checks are provided in Appendix K. Misleaders that require visual interpretation or contextual knowledge, such as misrepresentation or inappropriate axis ranges, are not covered by the linter. We evaluate its performance using both ground truth and predicted axis metadata.
Fine-tuned classifiers We train two classifiers: one that takes only the visualization image as input, and another that combines the image with its axis metadata. Visualizations are encoded using a frozen image encoder, while axis metadata is embedded using a frozen table encoder. We use the image encoder of TinyChart, a specialized chart understanding MLLM (Zhang et al., 2024), while the table encoder is TaPas (Herzig et al., 2020). For the second classifier, the resulting image embeddings are concatenated with the [CLS] token from the axis metadata embeddings. The image embedding or concatenated embeddings are passed through a trained multi-layer perceptron (MLP) classification head. The classifiers are trained on Misviz-synth to predict one misleader per visualization, or none.
Axis extraction The linter and the second classifier require axis metadata as input. While ground truth metadata is available for Misviz-synth, this is not the case for Misviz or real-world scenarios where only the visualization image is accessible. To address this, we implement an intermediate axis extraction step using DePlot (Liu et al., 2023). Since DePlot was originally only trained for chart-to-table extraction, we fine-tune it on (image, axis metadata) pairs from Misviz-synth.

### v2

@@ Abstract
Misleading visualizations are a potent driver of misinformation on social media and the web. By violating chart design principles, they distort data and lead readers to draw inaccurate conclusions. Prior work has shown that both humans and multimodal large language models (MLLMs) are frequently deceived by such visualizations. Automatically detecting misleading visualizations and identifying the specific design rules they violate could help protect readers and reduce the spread of misinformation. However, the training and evaluation of AI models has been limited by the absence of large, diverse, and openly available datasets. In this work, we introduce Misviz, a benchmark of 2,604 real-world visualizations annotated with 12 types of misleaders. To support model training, we also create Misviz-synth, a synthetic dataset of 57,665 visualizations generated using Matplotlib and based on real-world data tables. We perform a comprehensive evaluation on both datasets using state-of-the-art MLLMs, rule-based systems, and image-axis classifiers. Our results reveal that the task remains highly challenging. We release Misviz, Misviz-synth, and the accompanying code.111github.com/UKPLab/arxiv2025-misviz
@@ 1 Introduction
Misleading visualizations are charts that distort the underlying data, typically by violating design principles, leading readers to draw inaccurate conclusions (Tufte and Graves-Morris, 1983; Pandey et al., 2015; Lauer and O’Brien, 2020; McNutt et al., 2020; Lo et al., 2022; Lisnic et al., 2023; Lan and Liu, 2025). While many arise from unintentional design errors, misleading visualizations are also deliberately crafted by malicious actors to spread disinformation and manipulate public understanding, especially during crises such as the COVID-19 pandemic, where misleading charts circulated widely on social media (Correll and Heer, 2017; Tartaglione and de Wit, 2025; Lisnic et al., 2023). Prior work has shown that both humans (Pandey et al., 2014, 2015; O’Brien and Lauer, 2018; Yang et al., 2021; Ge et al., 2023; Rho et al., 2023) and MLLMs (Bendeck and Stasko, 2025; Chen et al., 2025; Pandey and Ottley, 2025; Tonglet et al., 2025) are easily deceived by such visualizations in question-answering tasks.
The deceptive features in these charts, or misleaders (Lisnic et al., 2023), often reside in subtle details easily missed by readers, such as axis tick intervals. Furthermore, misleaders are highly diverse: the latest taxonomies identify over 70 distinct types spanning a wide range of chart types, including bar charts, pie charts, and maps (Lo et al., 2022; Lan and Liu, 2025). In some cases, multiple misleaders affect the same visualization (Lo et al., 2022). Figure 1 shows 12 real-world examples.
@@
Automatically classifying whether a visualization is misleading and identifying which misleaders affect it, if any, can enable timely warnings to chart designers and readers and help prevent the spread of misinformation. This task is framed as a multi-label classification problem. While early work relied on rule-based systems called linters (Hopkins et al., 2020; Fan et al., 2022), recent studies have explored the use of MLLMs (Lo and Qu, 2025; Alexander et al., 2024). However, these approaches were evaluated on distinct datasets, which are either small or closed-access, limiting comparability and hindering progress.
In this work, we introduce Misviz, a large, diverse, and open benchmark comprising 2,604 real-world visualizations spanning 12 types of misleaders. It reflects scenarios in which detection models could flag visualizations published on the web. In Misviz, 70% of the visualizations contain up to three misleaders, while the remaining 30% are non-misleading. To support model training, we also release Misviz-synth, a synthetic dataset generated with Matplotlib (Hunter, 2007) using real-world data tables. Misviz-synth reflects scenarios in which detection models assist chart designers in identifying misleaders unintentionally introduced into their charts. The dataset includes not only the visualizations but also their underlying data tables, Python code snippets, and axis metadata, enabling the training of chart de-rendering models.
We conduct extensive experiments with three approaches: (a) state-of-the-art MLLMs, (b) a new rule-based linter that inspects axis metadata for design rule violations, and (c) new classifiers that take the visualization alone or in combination with the axis metadata as input. For (b) and (c), we fine-tune DePlot (Liu et al., 2023) to extract axis metadata as an intermediate step. Our experiments address the following research questions (RQs). RQ1: Which type of model performs best on real-world or synthetic instances? RQ2: Can detection models trained on synthetic instances generalize to real-world cases? RQ3: Can axis extraction models trained on synthetic instances generalize to real-world cases?
Our results show that MLLMs perform best on real-world visualizations, while linters and image-axis classifiers outperform them on synthetic ones, benefiting from the availability of training data for both axis extraction and misleader detection. While the fine-tuned DePlot can extract axes from Misvizsynth, it does not generalize well to Misviz, reducing the performance of the linter and classifier.
In summary, our contributions are as follows: (1) We introduce Misviz and Misviz-synth, the first large-scale open datasets for misleading visualization detection. (2) We propose a new linter and a new classification method that combines image and extracted axis metadata as input. (3) We conduct a comprehensive evaluation and error analysis, highlighting the strengths and weaknesses of each method and identifying directions for future work.
@@ 2 Related work
Dataset Instances Misleader types Chart types % non-misleading Open access Real-world Multi-label Axes, table, code
@@
Misvisfix (Das and Mueller, 5555) 450 74 > 5 20
Misviz-synth (ours) 57,665 12 5 39
@@
The first attempts to detect misleading visualizations relied on rule-based systems called linters (McNutt and Kindlmann, 2018; Hopkins et al., 2020; Chen et al., 2022). These linters assume the availability of the underlying data table or the chart code, which restricts their applicability to real-world scenarios. Fan et al. (2022) and Biselli et al. (2025) overcome these limitations by extracting tables using OCR tools from real-world visualizations before applying rule checks. However, the accuracy of these real-world linters depends heavily on the quality of the intermediate OCR step, which can vary widely (Biselli et al., 2025). Real-world linters have previously been evaluated in small-scale user studies with human-in-the-loop correction of OCR errors.
Others have explored the potential of MLLMs for the task. Lo and Qu (2025) evaluated four MLLMs with different prompts on a dataset of 150 real-world visualizations, sourced from the corpus of Lo et al. (2022) for misleading cases. They found that detection accuracy decreased as more misleader types were included in the prompt. Alexander et al. (2024) focused on GPT-4 (OpenAI, 2023), using visualizations from the social media platform X (Lisnic et al., 2023). However, access to this dataset requires a paid API, and reproducibility is further hindered by the platform’s frequent removal of posts. In a parallel work, Das and Mueller (5555) proposed a prompt with which SOTA MLLMs achieve high accuracy on a subset of the corpus of Lo et al. (2022).
@@
Table 1 compares prior datasets with Misviz and Misviz-synth. Misviz is over fifteen times larger than the dataset of Lo and Qu (2025). Unlike Alexander et al. (2024), it does not rely on paid APIs and ensures long-term access to all instances by archiving them on the Wayback Machine.222web.archive.org Misviz-synth is two to seven times larger than other synthetic datasets, and includes several more misleaders and chart types (Arif et al., 2024; Maciborski et al., 2025). In contrast to other synthetic datasets, Misviz-synth provides the underlying table, code, and axis metadata. The latter is necessary to fine-tune DePlot for axis extraction and answer our research questions.
@@ 3.1 Selected misleaders
Misviz covers 12 types of misleaders, selected from the 74 categories defined in the taxonomy of Lo et al. (2022), based on three key criteria. First, we excluded misleaders that are rarely observed in real-world scenarios. To determine this, we used misleader frequency statistics from the corpus of Lo et al. (2022) and discarded all categories with fewer than three instances. Second, we removed misleaders that are more likely to confuse than deceive the reader. As noted by Lo et al. (2022), the taxonomy includes both misleaders that distort the underlying data, the focus of this work, and others that may hinder readability or clarity without altering the interpretation of the data, such as missing titles or overplotting. Third, we excluded misleaders that require specific domain knowledge to identify. For example, using red to represent Democrats and blue to represent Republicans in a U.S. politics chart violates color conventions, but detecting this misleader requires familiarity with U.S. politics. Such misleaders require domain expertise that is not achievable with crowdworkers.
We define each selected misleader briefly below (Lo et al., 2022; Ge et al., 2023; Lan and Liu, 2025). They cover together 62.3% of all instances from the real-world corpus of Lo et al. (2022). Each of them is represented with an example in Figure 1 and in Appendix A.
Misrepresentation: the value labels displayed do not match the sizes of their visual encodings; e.g., bars may be drawn disproportionately to their corresponding numerical values.
@@
Inconsistent tick intervals: the ticks in one axis are evenly spaced, but their values are not, e.g., the tick values sequence is 10, 20, 40, 45.
@@ 3.3 Data labeling
3.4 Bounding box labeling
We recruited ten PhD students to annotate misleading visualizations using the VIA tool (Dutta et al., 2016; Dutta and Zisserman, 2019). Annotators drew bounding boxes around relevant misleader features, e.g., the initial tick mark on a truncated axis. Three misleaders were excluded: misrepresentation, 3D, and inappropriate use of pie chart, because bounding boxes are not suitable for representing them. Each student annotated up to 95 instances. We calculated IAA on a shared subset of 27 instances, three per misleader. Using an Intersection over Union (IoU) threshold of 0.4, the annotators achieved an IAA of 0.81, indicating strong agreement.
@@ 3.5 Data statistics
We did not leverage the large unlabeled corpus collected by Lo et al. (2022) to create Misviz. However, this resource could be used in future work to scale the dataset, as discussed in Appendix G.
@@ 4.1 Data collection
We collect real-world data tables from Our World in Data, an open-access data platform that covers a wide range of domains, including health, education, and the economy.555ourworldindata.org
@@ 4.2 Synthetic visualizations generation
Figure 2: The two-step process to create the synthetic visualizations of Misviz-synth based on real-world data.
@@
In the first step, given a raw real-world data table, we assign one or more types to each column. There are 10 different column types, explained in Appendix H. Based on these types, we select valid column combinations that require at least one numerical column and a natural key for each row. This natural key can consist of one or a combination of categorical or temporal columns. If the natural key involves multiple columns, we fix all but one to a constant value. For instance, if the numerical column is Win ratio and the natural key includes Football club and Year, we set Year to 2021 to compare the win ratios of football clubs for that year. Chart titles are automatically generated using templates based on the selected columns. We then determine which chart types, among bar, line, and pie charts, are suitable for the data. For example, line charts are only compatible with temporal data, not categorical data. Each valid configuration produces a “prepared” data table.
@@
For bar, line, and scatter plots, we automatically extract axis metadata. Since pie charts and maps lack axes, they do not have axis metadata. Axis metadata is stored as a table with four columns. (1) Seq indicates the sequential order of the tick marks along an axis. Horizontal axes are read from left to right, while vertical axes are read from bottom to top. (2) Axis indicates the name of the axis, e.g., x, y1, or y2, in case of dual axis. (3) Label is the tick label. (4) Relative position is a normalized float indicating the tick mark’s position, with spacing expressed relative to the distance between the first two tick marks. Examples of axis metadata are provided in Appendix I.
To increase visual diversity, we apply random variations in font size, background color, and axis-label positions. The complete list of variations is given in Appendix J.
Each instance is saved along with its title, data table, Matplotlib code, and axis metadata.
@@ 4.3 Data statistics
Misviz-synth contains 57,665 visualizations, distributed into stratified train-large (80.1%), train-small (9.6%), dev (3.1%), validation (3.2%), and test sets (4.1%). Misviz-synth covers the five most common chart types in Misviz: bar, line, and pie charts; scatterplots; and maps. Unlike Misviz, each visualization in Misviz-synth contains at most one chart type and one misleader, making it slightly less diverse than Misviz.
@@ 5.1 Baselines
Zero-shot MLLMs We evaluate the zero-shot capabilities on one run of the following commercial and open-weight MLLMs: GPT-4.1 and GPT-o3 (OpenAI, 2023), Gemini-2.5-Flash-Lite (Gemini-Team, 2024), Qwen-2.5-VL in 7B, 32B, and 72B variants (Qwen-Team, 2025), and InternVL3 in 8B, 38B, and 78B variants (Zhu et al., 2025). These models were selected for their strong performance on the ChartQA benchmark (Masry et al., 2022). The prompt includes the task description and definitions of the misleaders. The prompt is provided in Appendix K.
Rule-based linter We design a linter that detects misleaders by applying rule-based checks to the axis metadata of a visualization. Each misleader corresponds to a specific rule. If a visualization passes all checks, it is classified as having no misleaders. We rely on axis metadata rather than the underlying data table because it enables detection of a broader range of misleaders: truncated axis, inverted axis, dual axis, inconsistent tick intervals, inconsistent binning size, and inappropriate item order. Only the latter two misleaders could also be detected from the data table alone. Detailed descriptions of all rule checks are provided in Appendix L. Misleaders that require visual interpretation or contextual knowledge, such as misrepresentation or inappropriate axis ranges, are not covered by the linter. We evaluate its performance using both ground truth and predicted axis metadata.
Image-axis classifiers We train two classifiers: one that takes only the visualization image as input, and another that combines the image with its axis metadata. Visualizations are encoded using a frozen image encoder, while axis metadata is embedded using a frozen table encoder. We use the image encoder of TinyChart, a specialized chart understanding MLLM (Zhang et al., 2024), while the table encoder is TaPas (Herzig et al., 2020). For the second classifier, the resulting image embeddings are concatenated with the [CLS] token from the axis metadata embeddings. The image embedding or concatenated embeddings are passed through a trained classification head. The classifiers are trained on Misviz-synth to predict either one misleader per visualization or none.
Axis extraction The linter and the second classifier require axis metadata as input. While ground truth metadata is available for Misviz-synth, this is not the case for Misviz or real-world scenarios, where only the visualization image is accessible. To address this, we implement an intermediate axis extraction step using DePlot (Liu et al., 2023). Since DePlot was originally only trained for chart-to-table extraction, we fine-tune it on (image, axis metadata) pairs from Misviz-synth.